# CardioLive: Empowering Video Streaming with Online Cardiac Monitoring

### v1

@@ 6 Related Work
Cardiac Monitoring: Cardiac information is crucial for health monitoring, affective computing [68, 27] and deception analysis [17]. Traditional approaches in hospitals, e.g., electrocardiograms (ECGs) and CT scans [4], provide the most accurate data but require professional operation and are prohibitively expensive and cumbersome for everyday use. Recent advancements have focused on more portable solutions. Earable-based systems [18, 20, 28] allow earpieces to detect cardiac information, but they either need specific probing signals or custom hardware, limiting their widespread adoption. Similarly, wearable solutions necessitate constant wear, which is not practical for all users. Wireless technologies, including Wi-Fi [39], mmWave [70], and UWB [22], etc, are constrained by specific hardware which is not commonly available in video systems . Solutions using active acoustic sensing [58, 59, 48, 75] with smart speakers rely on pseudo-inaudible signals, which can be intrusive to human hearing and increase hardware burden. Video-based solutions use optical means to measure blood volume changes in tissues. Signal processing [23, 35, 61, 62] and deep learning [21, 40, 45, 34, 72, 73, 74, 41, 77] techniques have been developed to enhance these methods. Yet these solutions are sensitive to low light conditions, head/body movements, and typically perform poorly outside controlled environments. VocalHR [67] proves the potential of extracting heart rate from human speech. Although it leverages human speech effectively, it is limited by range, requires pre-calibration, and cannot distinguish multiple individuals. Differently, CardioLive is the first to combine the complementary and naturally co-existing audio and video modalities in online video streaming systems. Our video design incorporates temporal-frequency co-design and motion-aware aggregations for the first time in OCM to mitigate the light and body movement influence. The audio module employs the temporal acoustic filter for OCM. These designs are innovative and contribute to our performances.
@@
Video Streaming System: Video streaming systems have gained immense popularity due to their vast libraries of on-demand content, user-generated videos, and live streaming capabilities, catering to diverse viewer preferences, including YouTube, TikTok, Zoom, etc. They can be further categorized into VoD systems, live streaming systems and video conferencing systems. Research efforts have been devoted to communication protocols [30, 25], adaptive rate streaming algorithms [36, 64, 76], online learning [55, 29, 33, 71], etc. None of these works explore adding cardiac monitoring into modern video streaming systems. In contrast, CardioLive stands out as the first work that creates a middleware service of OCM that can be seamlessly integrated into mainstream video streaming systems.
7 Discussion and Future Work
@@
Audio-Video Pair: In our primary application scenarios (e.g., live streaming, online meetings, etc), audio and video naturally coexist. In practice, only video data is available in some situations, where CardioLive can be easily adapted to a video-only solution. Such periods can be detected through mature voice activity detection techniques [65]. Our results shown in Fig. 18 have demonstrated that CardioLive also performs well in video-only scenarios. CardioLive not only introduces a novel approach to OCM by utilizing audio-visual pairs for the first time, but also integrates these capabilities into a practical system with flexibility and robustness.
@@
Impacts on Original Streams: Integrating additional services into standard streaming platforms has been a bottleneck for many previous solutions [42, 19, 26]. In CardioLive, we address this challenge with a dedicated design of data hook and middleware service. Our approach ensures that these additional services are isolated from the original streams. With offscreen canvas, which operates independently in the extension, we avoid disrupting the original content. In meetings, our data hook duplicates data to the inference engine instantly, seamlessly, and without affecting the main video and audio streams. Our evaluations demonstrate that CardioLive operates without causing any disruptions or interference to ongoing streams.
@@
8 Conclusion
In this paper, we envision the attractiveness of Online Cardiac Monitoring (OCM) in video streaming and present CardioLive, the first-of-its-kind system to fuse both audio and video streams for online cardiac monitoring in video streaming systems. We devise an effective audio-visual network that can robustly and accurately unveil the nuanced cardiac activities, achieving an average MAE of 1.79 BPM and outperforming the video-only and audio-only solutions by 69.2% and 81.2%, respectively. Furthermore, we design and implement CardioLive as a plug-and-play microservice that can seamlessly be integrated into mainstream video streaming systems. We believe our work will significantly enhance the entertainment and healthcare value of video streaming and inspire a new direction in this field.

### v2

@@ VI Related Work
Cardiac Monitoring: Cardiac information is crucial for health monitoring, affective computing [yang2022survey, fairclough2020personal] and deception analysis [bian2024ubihr]. Compared with hospital solutions [HeartDiseaseSymptoms], recent advancements have focused on more portable solutions [xue2025ecg, alimbayeva2022portable, chan2025mobile, chan2019contactless]. Earable systems [cao2023heartprint, chen2024exploring, fan2023apg, cao2024earsteth, bui2019ebp, park2020heartquake] allow earpieces to detect cardiac information, but they either need specific probing signals or custom hardware, limiting their widespread adoption. Similarly, wearable solutions necessitate constant wear, which is not practical for all users. Wireless technologies, including Wi-Fi [liu2015tracking], mmWave [yang2016monitoring], and UWB [chen2021movi], etc, are constrained by specific hardware that is not commonly available in video systems. Solutions using active acoustic sensing [wang2023df, wang2022loear, qian2018acousticcardiogram, zhang2020your] with smart speakers rely on pseudo-inaudible signals, which can be intrusive to human hearing and increase hardware burden. Video-based solutions use optical means to measure blood volume changes in tissues. Signal processing [de2013robust, li2014remote, wang2016algorithmic, wang2015novel] and deep learning [chen2018deepphys, liu2020multi, niu2020video, li2023learning, yu2019remote, yu2019remoteCompress, yu2023physformer++, liu2023efficientphys, zou2024rhythmformer, qian2024cluster, wang2024rppg, zou2025rhythmmamba, luo2024physmamba, wu2025cardiacmamba] techniques have been developed to enhance these methods. Yet these solutions are sensitive to low light conditions, head/body movements, and typically perform poorly outside controlled environments. VocalHR [xu2022hearing] proves the potential of extracting heart rate from human speech. However, it is limited by range and requires pre-calibration. Differently, CardioLive is the first to combine the complementary and naturally co-existing audio and video modalities in online video streaming systems. Our video design incorporates temporal-frequency co-design and motion-aware aggregations for the first time in OCM to mitigate the light and body movement influence. The audio module employs the temporal acoustic filter for OCM. These designs are innovative and contribute to our performances.
@@
Video Streaming System: Video streaming systems have gained immense popularity due to their vast libraries of on-demand content, user-generated videos, and live streaming capabilities, catering to diverse viewer preferences, including YouTube, TikTok, Zoom, etc. They can be further categorized into VoD systems, live streaming systems and video conferencing systems. Research efforts have been devoted to communication protocols [hamadanian2023ekho, dhawaskar2023converge], adaptive rate streaming algorithms [li2023dashlet, wen2023adaptivenet, zhou2019learning, li2022livenet, tashtarian2024artemis], online learning [tang2023lut, guan2023metastream, khani2023recl, yi2023boosting, zhang2024aralive, wu2024toward, chen2024videollm, qian2024streaming], and video understanding and serving [li2025videoscan, tong2022videomae, naiman2025lv, du2020server], etc. None of these works explores adding cardiac monitoring to modern video streaming systems. In contrast, CardioLive stands out as the first work that creates a middleware service of OCM that can be seamlessly integrated into mainstream video streaming systems.
VII Discussion and Future Work
@@
Audio-Video Pair: In our primary application scenarios (e.g., live streaming, online meetings, etc), audio and video naturally coexist. In practice, only video data is available in some situations, where CardioLive can be easily adapted to a video-only solution. Such periods can be detected through mature voice activity detection techniques [wiseman2016python]. Our results shown in Fig. 19 have demonstrated that CardioLive also performs well in video-only scenarios. CardioLive not only introduces a novel approach to OCM by utilizing audio-visual pairs for the first time, but also integrates these capabilities into a practical system with flexibility and robustness.
@@
Impacts on Original Streams: Integrating additional services into streaming platforms can be a bottleneck for many previous solutions [liu2020grad, mmdetection, du2020server]. In CardioLive, we address it with a dedicated design of data hook and middleware service. Our approach ensures that these additional services are isolated from the original streams. With an offscreen canvas, we avoid disrupting the original content. In meetings, our data hook duplicates data to the inference engine instantly without affecting the main video and audio streams. Our evaluations demonstrate that CardioLive operates without causing any disruptions or interference to ongoing streams.
@@
VIII Conclusion
In this paper, we envision the attractiveness of Online Cardiac Monitoring (OCM) in video streaming and present CardioLive, the first system to fuse both audio and video streams for OCM. We devise an effective audio-visual network that can robustly and accurately unveil the nuanced cardiac activities, achieving an average MAE of 1.79 BPM and outperforming the video-only and audio-only solutions by 69.2% and 81.2%, respectively. Furthermore, we design and implement CardioLive as a plug-and-play middleware that can seamlessly be integrated into mainstream streaming systems. We believe our work will significantly enhance the entertainment and healthcare value of video streaming and inspire new directions.
Acknowledgments
This work is supported by the NSFC under Grant No. 62222216, the Hong Kong RGC ECS under Grant No. 27204522, and GRF under Grant No. 17212224.